please help me solve this quadratic equation in vertex form.
y=x^2+12x-4
thank you very much

Answer by nerdybill(7384)

The vertex form is:
y= a(x-h)^2+k
.
This involves "completing the square".
For reference see:
http://www.purplemath.com/modules/sqrvertx.htm
.
y=x^2+12x-4
y+4 = x^2+12x
y+4+36 = x^2+12x+36
y+40 = (x+6)^2
y = (x+6)^2 - 40
y = (x-(-6))^2 + (-40)
.
Now, we see that:
(h,k) = (-6,-40)